Commissioning Notes for 20080418 UT

Observers: J. Hill, M. Wagner plus T. Sargent, P. Grenz, M. Gusick at the beginning. and D. Miller for the second half
Telescope Operator: A. Ceranski

Summary

The wind has calmed enough to open, although there is still a pretty stiff breeze. Lost 1 hour to a pinched fiber in the rotator electronics. Lost 1 hour to TCS being in a confused state. Spent an hour chasing various issues with irc/irs. Had significant problems with socket timeout errors between irs and irtc. Spent about half the night collecting cubes of data to analyze with EF. Mark and Doug haven't yet succeeded in making EF produce numbers that we believe. We spent the last couple hours teaching Doug everything we know. He too can now compress a data cube into a single pupil and collimate the telescope by eye.

The plan for tomorrow night is to have Doug fire up the AGw and take some images with it. This will allow the first version of GCS to have the correct numbers for the basic parameters. John and Mark are going down, but John will spend a couple hours tomorrow observing from Tucson. If the AGw works OK in engineering mode, then we can go back to trying to close the loop in EF.

Nighttime Operations

01:35 Open dome facing east before sunset, T=-1degC, Photometric sky, Wind a mere 10 m/sec and remarkably steady.

Rotator Problems

01:40 Rotator won't come on. SERDES Framing Error. Fortunately we have Mike who designed the circuit right here in the control room, so debugging is quite efficient. Otherwise we might have spent half the night chasing this one.

02:14 takeseq 2 sec x 15 at zenith, ~7000 counts/pix

First try had a socket error on the single image. IRTC took the images, but Dave's takeseq script reported failure. "irc GetSeqImages 15" reports failure even though we seem to be getting the image OK on the Repository disk. irs reports [E] Exception was caught:[Error] IRSocket::GetSequenceImages -> Could not read from socket.

02:50 A temporary fiber link between upper treehouses was crushed in the hatch. We ran a temporary replacement for the temporary fiber. Restarted MCSPU, restarted PCS and everybody is happy.

TCS Problems

03:00 Found that netconfig was active and PSFL, PSFR, ECS were running on to-station. Stopped them, started passive on to-station and restarted the subsystems.

Trying takeEFpair

08:21 irc Move 0. 0. 0. 0. 0. 0. 0 ABS M1 0 left > Problem with move command. res=1

...Done! (failed: see LOGs/takeEFpair_032116.log)

irs says: [I] Move command status: FAILED [I] Invalid buffer:

Stopped and restarted CSQ. Telescope stops per plan.

pointto WT10_399 again [I] PresetTelescope command status: FAILED [I] PCS set new target to RA: 2.916283373516454e+00 Dec: 1.255920361371133e+00 complete [I] Invalid buffer: [I] Invalid buffer: [I] -21

Restarted all of TCS and the invalid buffers are gone.

IRTC Problems

04:22 [E] Exception was caught:[Error] IRSocket::GetImage::Header -> Time out: Could not read from socket. (multiple times)

04:30 Reboot the irtc windows machine (Using Crtl-Alt-Del to bring up task manager.) This made the socket interface better, but it is still having problems.

GetSeq fails from irs every time.

Collecting some intra/extra/sky pairs on sky

04:56 Successfully collected first intra/extra/sky pair irtc2_20080417.285634.fits a_00005_cube.fit irtc2_20080417.285719.fits a_00006_cube.fit irtc2_20080417.285804.fits a_00007_cube.fit in001.fits ex001.fits

05:42 Notice that Position angle is different than it was before we restarted TCS. Now it is rotated 45 deg from last night even though it was OK at the beginning of tonight. Stopped and restarted PCS. Then PA is OK again. I am suspicious that PCS is not rereading PCSInstrument.conf on a change of authorization.

Talked to Jose on the phone. Jose thinks the GetSequence problem is because he was using the frame rate to estimate the time to wait, but this is not the dominant parameter when exposures are longer than 1/15 second. He doesn't have any solution for the socket problem. He is going to make a simulator to exercise the irs socket library.

06:08 WT10_399 is now at offset -10,0 with SX M1 Global Offsets of X=0 Y=-0.5 Z=-0.21. (Implies that my Y temperature coefficient has the wrong sign?)

We are using inex_seq.cl since it doesn't have the error checking that takeEFpair.cl does,

06:34 Our next problem is that the move command works OK in moving X,Y or Z, but it fails for moving the angles. Fails in this case means reporting success even though the secondary does not move in angle.

irc Move 1. 0. 0. 0. 1. 0. 17 ABS M2 0 left

The above example moves X=1 successfully, but does not get any value of RY into PSF. I have tried both arcsec and degrees as the units.

Some good news is that our open loop tracking is quite nice. The image is stable for several minutes.

06:45 Another cl < inex_seq.cl, but again we have problems with socket errors so we missed the first single image. (14,15,16)

Note that inex_seq.cl has 10 sec sleep commands in it to give the M2 time to finish moving as Paul hasn't yet update the command return in OSS.

06:55 pointto WT10_354 And it landed IN THE FOV!, it needed an offset of only 3,3

07:16 cl < inex_seq.cl at EL=70.9 This one has X=+1 mm instrument offset. (17,18,19) LM1X=1.34

07:32 cl < inex_seq.cl at EL=72.5 This one has X=0 mm instrument offset. (20,21,22) LM1X=0.30

Doug M. arrives from Tucson.

07:47 cl < inex_seq.cl at EL=72.5 This one has X=0 mm instrument offset. (23,24,25)

offset -4 7

07:53 cl < inex_seq.cl at EL=74.5 This one has X=0 mm instrument offset. (26,27,28)

offset -7 7

08:04 cl < inex_seq.cl at EL=75.5 These are cubes of 30 images to get more averaging. (29,30,31)

Opened vent doors with wind about 5 m/sec.

08:20 pointo WT10_396, star lands on top edge of field.

08:30 Aaron heads off to the utility building to restart the chiller. I've stopped mirror ventilation to reduce the load on the system.

08:40 Verfied that the LBTtools.IRTC.focus script works correctly in both ABS and REL mode. Also verified that the first command of inex_seq.cl does not overwrite this because it uses DFLAG=0.

08:46 cl < inex_seq.cl at EL=51.3 These are cubes of 30 images to get more averaging. (32,33,34) Yg=-1.1 to give rough coma correction. LM1Y=-1.35

offset -3 30 after changing collimation

08:57 cl < inex_seq.cl at EL=52.0 These are cubes of 30 images to get more averaging. (35,36,37) Yg=-0.1 to add coma. LM1Y=-0.36

09:05 At Rotator Angle = 121, Azimuth=17, Parallactic=-14 on WT10_396 the command
irc OffsetGuiding 0.0 -5.0 0.0 RADEC_FOCAL left
Moves the image toward the lower right of the image.
While the offset -5 0 command moves the image straight to the right.

Mark and Doug have not yet been able to match the amplitude of coma that I've been applying by moving the mirror in X,Y. There is still a factor of three missing.

09:23 Start the training exercise for Doug to learn to start all the systems.

10:11 Star WT10_345 at top of field for PA=0, at left of field for PA=90, doesn't respond to PA=270, at right of field for PA=-90.

10:34 cl < inex_seq.cl at EL=73 These are cubes of 30 images to get more averaging. (EF/2,3,4)

10:35 Restart mirror ventilation

Trying to collimate by eye to see if the LBC ByEyeAO directions are still applicable. (Apparently yes, Coma is OK, Astig Z6 and Spherical Z11 are OK, we still need to check for mirror reflections in other aberrations.)

Really bright pupils

11:50 pointto M5_0699. It landed on the upper left corner of the chip.

Seeing monitor says 1 arcsec seeing, but pupils are dominated by a lot of slow moving bubbles that look like dome seeing. Note that while these cubes have very high S/N, they will not be useful for EF active optics analysis since the 30 frames are taken very close together in time. This rapid movie give insufficient averaging of the atmosphere.

12:20 cl < inex_seq.cl with 50 ms exposures. b__00006,7,8, but some pupils are saturated.

12:26 repeat with 2 ms exposures, but still parts are saturated. b__00009,10,11

12:28 repeat with 0.5 ms exposures. b__00012,13,14

12:30 Close dome. Wind has dropped to dead calm in the last 20 minutes.

-- JohnHill - 18 Apr 2008
Topic revision: r6 - 19 Apr 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback